OpenStudy (anonymous):

7x − 2(x + 1) = 6x + 14 −12 12 -16 16

OpenStudy (anonymous):

7x - 2x +1 = 6x+14

OpenStudy (anonymous):

Now get x on one side.

OpenStudy (anonymous):

7x-2x-6x=-1+14

OpenStudy (amistre64):

you distributed incorrectly to start with

OpenStudy (nurali):

7x − 2(x + 1) = 6x + 14 7x − 2x - 2= 6x + 14 5x-2=6x+14 6x-5x=-2-14 x=-16

OpenStudy (anonymous):

answeer?

OpenStudy (anonymous):

lol i did didn't I...

OpenStudy (anonymous):

The answer is -16... beaause x = -16

OpenStudy (amistre64):

we would prefer you participate in the solution process as opposed to simply asking for an answer

OpenStudy (anonymous):

Yeah because if we simply give the answer it's a form of cheating.

OpenStudy (uri):

7x-2x-2-6x+14 5x-2=6x+14 -2-14=6x-5x -16=x or x=-16